# Generalized Hypercube Queuing Models with Overlapping Service Regions

### v2

@@ 2.2.4 Balance equations
However, the above balance equations cannot be solved computationally as the number of states on the hyperlattice is infinite and grows exponentially with the number of servers even when queue capacity is limited. To overcome this difficulty, we introduce an efficient estimation strategy in the next section.
2.3 Efficient steady-state distribution estimation
Now we introduce an efficient computational method to estimate the steady-state distribution that particularly leverages the symmetrical structure of the state space and the sparsity of its transition rate matrix.
2.3.1 Aggregated states and queue length distribution
We first exploit the fact that the hyperlattice queue model in the aggregate is simply a birth-death model whose states can be represented by $\{N_{k}\}_{k=0,1,\dots,\infty}$ and $k$ can be interpreted as the number of calls currently in the system. This equivalence is obtained by collecting together all states having an equal value of $k=\left\lVert B\right\rVert_{1}$ as shown in Figure 8 (b). Their summed steady-state probability is equal to the comparable probability of state occurring in the birth-death model, i.e.,
$\sum_{B:\left\lVert B\right\rVert_{1}=k}\mathbb{P}\{B\}=\mathbb{P}\{N_{k}\}.$ (5)
The birth and death rates of the aggregated model can be defined by the following proposition:
Proposition 2.
The birth rate and death rate of the aggregated state $N_{k}$, represented by $\bar{\lambda}_{k}$ and $\bar{\mu}_{k}$ respectively, can be expressed as follows:
$\bar{\lambda}_{k}=$ $~{}\binom{k+I-1}{I-1}\lambda,\quad k=0,1,\dots,$ (6)
$\bar{\mu}_{k}=$ $~{}\binom{k+I-2}{I-1}\mu,\quad k=1,2,\dots.$
Proof.
To begin with, we define $\bar{\lambda}\left(B_{u}\right):=\sum_{i\in\mathcal{I}}q_{uu^{i+}}$ as the total sum of upward transition rates originating from an arbitrary state $B_{u}$. Given the upward transition rates that leave state $B_{u}$ defined in (1), we have
$\bar{\lambda}\left(B_{u}\right)$ $=\sum_{i\in\mathcal{I}}q_{uu^{i+}}=\sum_{i\in\mathcal{I}}\left[\lambda_{i}+\sum_{e\in\mathcal{E}}\eta_{e,u}(i)\lambda_{e}\right]=\sum_{i\in\mathcal{I}}\lambda_{i}+\sum_{e\in\mathcal{E}}\lambda_{e}\left[\sum_{i\in\mathcal{I}}\eta_{e,u}(i)\right]$
=$\stackrel{{\scriptstyle(i)}}{{=}}\sum_{i\in\mathcal{I}}\lambda_{i}+\sum_{e\in\mathcal{E}}\lambda_{e}=\lambda$
where the equation $(i)$ holds due to (3). Hence, the sum of all upward transition rates leaving any given state is constantly $\lambda$.
Now the birth rate of the aggregated state $N_{k},~{}k\geq 0$ can be obtained by multiplying $\bar{\lambda}(B_{u})$ by the number of states that are part of $N_{k}$, which is equivalent to $\binom{k+I-1}{I-1}$. It is also evident that the total of all downward transition rates leading to an arbitrary state is $\sum_{i\in\mathcal{I}}\mu_{i}$, which is equal to $\mu$. Consequently, we can obtain the death rate of $N_{k},~{}k\geq 1$ similarly by multiplying the total service rate $\mu$ by the number of states included in $N_{k}$. ∎
Remark 1.
The birth and death rates in the aggregated model are solely determined by the total arrival rate $\lambda$ and service rate $\mu$, and both rates increase exponentially with respect to $k$. If we let $n=k+I-1$ and $m=I-1$, we can use Stirling’s Formula [18] to obtain the asymptotic expression for $\binom{n}{m}$:
$\frac{\sqrt{2\pi n}(n/e)^{n}}{m!\sqrt{2\pi(n-m)}((n-m)/e)^{n-m}}=(1-m/n)^{m-1/2}\frac{e^{-m}}{m!(1-m/n)^{n}}n^{m}.$
It is worth noting that the first term on the right-hand side of the equation approaches 1 as $n\rightarrow\infty$. Additionally, the denominator of the middle term on the right-hand side approaches $m!e^{-m}$, which means that the middle term also approaches $1/m!$. Therefore, we can approximate the rate of growth as follows when $k\rightarrow\infty$:
$\binom{k+I-1}{I-1}\approx\frac{(k+I-1)^{I-1}}{(I-1)!}.$
Figure 11 shows the effect of $I$ and $k$ on the birth and death rates of hyperlattice queuing models.
(a) Two servers
(b) Three servers
(c) Four servers
Figure 11: The effect of the number of servers on the birth and death rates of hyperlattice queuing models with $\lambda=\mu=I$. The upward transition rates, both from non-saturated states (represented by gray dashed line) in $N_{k}$ and saturated states (represented by red dashed line) in $N_{k}$, to states in $N_{k+1}$ are shown collectively.
Proposition 3 (Distribution of aggregated queue length).
The probabilities of aggregated states can be written in closed-form using (6) as in a birth-death model
$\mathbb{P}\{N_{k}\}=\left(\prod_{l=1}^{k}\frac{\bar{\lambda}_{l-1}}{\bar{\mu}_{l}}\right)\left(\sum_{k^{\prime}=1}^{\infty}\frac{\prod_{l=0}^{k^{\prime}-1}\bar{\lambda}_{l}}{\prod_{l=1}^{k^{\prime}}\bar{\mu}_{l}}\right)^{-1}.$ (7)
Proof.
The proof is as follows. In a steady state for a continuous-time Markov chain, the rate at which individuals transition into a particular state must be equal to the rate at which individuals transition out of that state. Recall that the birth and death rate of aggregated state $N_{k}$ are denoted by $\bar{\lambda}_{k}$ and $\bar{\mu}_{k}$, respectively, and $\mathbb{P}\{N_{k}\}$ is the probability of being in state $N_{k}$, then we have the set of equations:
$\bar{\lambda}_{0}\mathbb{P}\{N_{0}\}=$ $~{}\bar{\mu}_{1}\mathbb{P}\{N_{1}\},$
$(\bar{\lambda}_{k}+\bar{\mu}_{k})\mathbb{P}\{N_{k}\}=$ $~{}\bar{\lambda}_{k-1}\mathbb{P}\{N_{k-1}\}+\bar{\mu}_{k+1}\mathbb{P}\{N_{k+1}\},\quad k>1.$
The first equation gives us $\mathbb{P}\{N_{1}\}=(\bar{\lambda}_{0}/\bar{\mu}_{1})\mathbb{P}\{N_{0}\}$ and the second equation suggests that:
$\mathbb{P}\{N_{k+1}\}=\frac{(\bar{\lambda}_{k}+\bar{\mu}_{k})\mathbb{P}\{N_{k}\}-\bar{\lambda}_{k-1}\mathbb{P}\{N_{k-1}\}}{\bar{\mu}_{k+1}}=\frac{(\bar{\lambda}_{k}+\bar{\mu}_{k})\mathbb{P}\{N_{k}\}-\bar{\mu}_{k}\mathbb{P}\{N_{k}\}}{\bar{\mu}_{k+1}}=\frac{\bar{\lambda}_{k}}{\bar{\mu}_{k+1}}\mathbb{P}\{N_{k}\}.$
Therefore, we have
$\mathbb{P}\{N_{k}\}=\frac{\prod_{l=0}^{k-1}\bar{\lambda}_{l}}{\prod_{l=1}^{k}\bar{\mu}_{l}}\mathbb{P}\{N_{0}\}.$ (8)
Assuming that the birth and death rates are such that we actually have well-defined steady-state probabilities, $\sum_{k=0}^{\infty}\mathbb{P}\{N_{k}\}=1$, and so
$\sum_{k=1}^{\infty}\frac{\prod_{l=0}^{k-1}\bar{\lambda}_{l}}{\prod_{l=1}^{k}\bar{\mu}_{l}}\mathbb{P}\{N_{0}\}=1,$
which implies the desired result. ∎
The probabilities of the “tail” of the hyperlattice (states with large $k$ value) are usually low and can be negligible since their chance of happening decreases with the length of the queue. As a result, in practice, we can ignore the tail of the hyperlattice and focus on the states where the total number of calls in the system is less than $K$, or $\left\lVert B\right\rVert_{1}\leq K$. We represent this set of states by $\mathcal{U}_{K}=\{0,\dots,U_{K}\}$, where $U_{K}$ denotes the size of the set.
Lemma 1.
@@ 2.3.1 Aggregated states and queue length distribution
Proof.
The size of the set $\mathcal{U}_{K}$ is equal to the summation $U_{K}=\sum_{k=0}^{K}\binom{k+I-1}{I-1}$, which can be simplified using the Hockey Stick identity [12]:
@@
Applying the identity with $m=I-1$ and $n=K+I-1$, we get:
@@
Therefore, the result of the sum is $\binom{K+I}{I}$. ∎
2.3.2 Computing steady-state distribution
@@
where the right-hand side is a column vector of zeros. By using equation (5), the sum of probabilities in set $\mathcal{U}_{K}$ denoted by $\sigma_{K}$ is equal to the sum of probabilities in set $\{N_{k}\}$ up to $K$:
$\sigma_{K}=\sum_{u\in\mathcal{U}_{K}}\mathbb{P}\{B_{u}\}=\sum_{k=0}^{K}\mathbb{P}\{N_{k}\}.$ (10)
Therefore, we can estimate $\pi_{K}$ by solving (9) and (10). However, there is redundancy in these equations as the number of equations is $K+1$ and the number of unknown variables is $K$. To eliminate this redundancy, we replace one row of $Q_{K}$ with a row of ones, usually the last row, and denote the modified matrix as $Q^{\prime}_{K}$. Similarly, we modify the “solution” vector, which was all zeros, to be a column vector with $\sigma_{K}$ in the last row and zeros elsewhere, and denote it as $e_{K}$. In practice, we solve the resulting equation:
@@
Figure 12: Support of the transition rate matrix $Q_{K}$ of hyperlattice queuing models (without diagonal entries). The shaded entries correspond to the matrices’ non-zero entries and the white spaces are zero.
@@
We note that the solution to the above balance equations requires a matrix inversion of $Q_{K}$ (or $Q^{\prime}_{K}$), which can become computationally challenging for large matrix sizes. By Lemma 1, the size of $Q_{K}$ can reach a staggering $U_{K}\approx(K+I)^{I}/I!$ (using Stirling’s approximation [18]) even for moderate values of $K$ when the system has $I>2$ servers, presenting a considerable computational hurdle. Fortunately, the sparsity of transition rate matrices (as illustrated in Figure 12) can be leveraged to iteratively find the steady-state distribution using the power method [16], by constructing a probability transition matrix of a discrete-time Markov chain through uniformization of the original continuous-time Markov chain. Let $\mathbb{I}$ represent the identity matrix and denote the maximum absolute value on the diagonal of the transition rate matrix $Q_{K}$ as $\gamma=\max|q_{uu}|$. By initializing with a vector $\pi_{K}^{(0)}$, we can compute $\pi_{K}^{(t)}=\pi_{K}^{(t-1)}(\mathbb{I}+Q_{K}/\gamma)$ at each iteration $t=1,2,\dots$, until the distance between $\pi_{K}^{(t)}$ and $\pi_{K}^{(t-1)}$ falls below a specified threshold. The convergence of this approach is guaranteed, and the steady-state distribution can be efficiently computed despite the potentially large size of the transition rate matrix.
@@ 2.4 Measures of performance
In this section, we explicate two key performance measures for a service system with overlapping service regions, using the proposed hyperlattice queuing model as a means of analysis.
@@ 2.4.1 Individual workloads
The individual workloads can be further used to calculate various types of system-wide workload imbalance defined in [13].
@@ 2.4.2 Fraction of dispatches
In addition, we use $\rho_{i,i}$ to represent the fraction of dispatches that send the server $i$ to its primary service region $i$, which is consistently equal to $\lambda_{i}/\lambda$ since this region can exclusively be handled by server $i$. We note that $\rho_{i,e}=0$ if $i\notin e$ and $\sum_{i,k}\rho_{i,e}+\sum_{i}\rho_{i,i}=1$. Additionally, in practice, we consider $u\in\mathcal{U}_{K}$ rather than $\mathcal{U}$ in (12) as we only calculate the steady-state probabilities for the first $U_{K}$ states according to Section 2.3.
Remark 2.
The proportion of dispatches can be employed to derive expressions for travel time metrics. As per Larson (1974), these metrics can be obtained similarly. It is important to note that this is based on the assumption that queuing delay will be the predominant factor affecting the server’s response time for each incident.

### v4

@@ 3.2.6 Balance equations
However, one can observe that the above balance equations are hard to solve as the number of states on the hyperlattice is infinite and grows exponentially with the number of servers, even when the queue capacity is limited. Additionally, from Equation (4), the transition rates $q_{u,v}$ explicitly depend on the state-dependent dispatch policy $\eta_{e,u}$ for all $e\in\mathcal{E}$ and $u,v\in\mathcal{U}$. This state-dependent structure further complicates solving Equations (6), rendering them computationally intractable. To overcome this difficulty, we introduce an efficient approximation through a truncated hyperlattice model in the following section.
3.3 A truncated hyperlattice model
We approximate the solution to Equations (6) by solving the balance equations for a truncated hyperlattice model, in which the total capacity across all servers is bounded by an upper bound $K$. Specifically, instead of modeling each server as an infinite-capacity queue, we model the entire system analogously to an aggregated Erlang loss queue, i.e., the arrival rate in any region is set to zero whenever the total number of units in the system reaches $K$. In a light-traffic regime, setting the upper bound $K$ sufficiently large allows the truncated hyperlattice model to approximate the stationary dynamics of the original infinite-capacity queues closely. Figure 3.2.1(b) illustrates an example of a 2-server overlapping queueing system truncated at $K=2$, with the original system shown in Figure 3.2.1(a).
3.3.1 Approximate stationary distribution under truncation
Rather than attempting to calculate the stationary distribution across an infinite number of states in the original state space, the problem can be simplified by focusing only on a finite subset of states. These states are selectively considered if their $k$ values are less than a user-defined hyper-parameter $K$, or $\left\lVert B\right\rVert_{1}\leq K$, and the sum of their steady-state probabilities remains to be one. It is worth noting that the stationary probability of the “tail” of the hyperlattice (states with large $k$ value) is usually low and can be negligible since their chance of happening decreases with the length of the queue. Specifically, in a light-traffic regime (e.g., when the arrival rate is significantly smaller than the service rate), the steady-state probability in the tail of the hyperlattice decreases fast as $K$ increases. As a result, the truncated queueing system provides a close approximation of the original system when $K$ is sufficiently large, where the stationary tail probability is negligible in the original system. In contrast, in a high-traffic regime, the truncated hyperlattice model may diverge significantly from the original model with infinite capacity. However, for certain performance metrics, e.g., workload variance across servers, the truncated model can still serve as a reasonable approximation of the original model. The accuracy of this approximation in a high-traffic regime depends primarily on how sensitively the metric of interest responds to the system state. In the following experiments, we demonstrate that when focusing on server workload, the truncated model remains a good approximation of the simulation results.
We represent this set of truncated states by $\mathcal{U}_{K}=\{0,\dots,U_{K}\}$, where $U_{K}$ denotes the size of the set. In the subsequent discussion, Lemma 3.3 presents the equation for $U_{K}$ in relation to $K$ and $I$. This demonstrates that the size of the set can expand exponentially relative to the number of servers $I$, and it can become quite substantial even for moderate values of $K$, particularly in systems with multiple servers.
Lemma 3.3
@@ 3.3.1 Approximate stationary distribution under truncation
Proof 3.4
Proof of Lemma 3.3 The size of the set $\mathcal{U}_{K}$ is equal to the summation $U_{K}=\sum_{k=0}^{K}\binom{k+I-1}{I-1}$, which can be simplified using the Hockey Stick identity Jones (1996):
@@
Applying the identity with m=I1m=I-1 and n=K+I1n=K+I-1, we get:
@@
Therefore, the result of the sum is $\binom{K+I}{I}$.
@@
where the right-hand side is a column vector of zeros. The sum of probabilities in set $\mathcal{U}_{K}$ is equal to 1, i.e.,
$\sum_{u\in\mathcal{U}_{K}}\mathbb{P}\{B_{u}\}=1.$ (8)
Therefore, we can compute $\pi_{K}$ by solving (7) and (8). However, there is redundancy in these equations as the number of equations is $K+1$ and the number of unknown variables is $K$. To eliminate this redundancy, we replace one row of $Q_{K}$ with a row of ones, usually the last row, and denote the modified matrix as $Q^{\prime}_{K}$. Similarly, we modify the “solution” vector, which was all zeros, to be a column vector with one in the last row and zeros elsewhere, and denote it as $e_{K}$. In practice, we solve the resulting equation:
@@
Support of the transition rate matrix $Q_{K}$ of hyperlattice queueing models (without diagonal entries). The shaded entries correspond to the matrices’ non-zero entries, and the white spaces are zero.
@@
We note that the solution to the above balance equations requires a matrix inversion of $Q_{K}$ (or $Q^{\prime}_{K}$), which can become computationally challenging for large matrix sizes. By Lemma 3.3, the size of $Q_{K}$ can reach a staggering $U_{K}\approx(K+I)^{I}/I!$ (using Stirling’s approximation (Romik 2000)) even for moderate values of $K$ when the system has $I>2$ servers, presenting a considerable computational hurdle. Fortunately, the sparsity of transition rate matrices (as illustrated in Figure 3.3.1) can be leveraged to iteratively find the steady-state distribution using the power method (Nesterov and Nemirovski 2015), by constructing a probability transition matrix of a discrete-time Markov chain through uniformization of the original continuous-time Markov chain. Let $\mathbb{I}$ represent the identity matrix and denote the maximum absolute value on the diagonal of the transition rate matrix $Q_{K}$ as $\gamma=\max|q_{uu}|$. By initializing with a vector $\pi_{K}^{(0)}$, we can compute $\pi_{K}^{(t)}=\pi_{K}^{(t-1)}(\mathbb{I}+Q_{K}/\gamma)$ at each iteration $t=1,2,\dots$, until the distance between $\pi_{K}^{(t)}$ and $\pi_{K}^{(t-1)}$ falls below a specified threshold. The convergence of this approach is guaranteed, and the steady-state distribution can be efficiently computed despite the potentially large size of the transition rate matrix.
3.3.2 Analysis of two canonical policies
Although the stationary distribution of a general state-dependent policy is often intractable to derive, two canonical policies admit explicit characterizations: the Static Random-Routing policy and the Join-the-Idle-Queue policy. We analyze these policies under the truncated hyperlattice model and show that the truncated approximation provides a valid method in the light-traffic regime.
Static Random-Routing policy.
Consider a state-independent static random-routing policy $\eta$. Specifically, recall that $\eta_{e}$ is a vector denoting the probability distribution over servers that determines how an arrival from region $e$ is routed. Then, the system is equivalent to $|\mathcal{I}|$ independent $M/M/1$ queues, where the arrival rate to server $i$ is $\lambda_{i}+\sum_{e\in\mathcal{E}}\eta_{e}(i)\lambda_{e}$ and the service rate is $\mu_{i}$. Assume the stability condition holds, i.e., $\lambda_{i}+\sum_{e\in\mathcal{E}}\eta_{e}(i)\lambda_{e}<\mu_{i}-\delta_{i}$ for each $i\in\mathcal{I}$ and some $\delta_{i}>0$. Then, for server $i$, the stationary distribution of the number-in-system is the geometric distribution $\pi_{i}(n)=(1-\rho_{i})\rho_{i}^{n}$, where $\rho_{i}=\frac{\lambda_{i}+\sum_{e\in\mathcal{E}}\eta_{e}(i)\lambda_{e}}{\mu_{i}}$ is the traffic intensity at server $i$. Then, let $N_{i}$ be a random variable with distribution $\pi_{i}$, and define $S:=\sum_{i\in\mathcal{I}}N_{i}$ as the total number-in-system across all servers. The stationary probability of the truncated state space $\mathcal{U}_{K}$ is $\mathbb{P}\!\left\{S\leq K\right\}$. The probability generating function of $N_{i}$ is
$G_{i}(z)=\mathbb{E}\left[z^{N_{i}}\right]=\frac{1-\rho_{i}}{1-\rho_{i}z},\quad 0\leq z<\frac{1}{\rho_{i}}.$
Since each server is independent, the probability generating function of $S$ is $\prod_{i\in\mathcal{I}}G_{i}(z)$. Then, for any $z\in\left(1,1/\rho_{\max}\right)$ with $\rho_{\max}:=\max_{i}\rho_{i}$, Markov’s inequality on $z^{S}$ yields
$\mathbb{P}\{S\geq K\}=\mathbb{P}\left\{z^{S}\geq z^{K}\right\}\leq z^{-K}\prod_{i\in\mathcal{I}}\frac{1-\rho_{i}}{1-\rho_{i}z}\leq C(z)e^{-K\ln z},$
where $C(z):=\prod_{i\in\mathcal{I}}\frac{1-\rho_{i}}{1-\rho_{i}z}$ is a finite constant given $z$. Then, for any $z\in\left(1,1/\rho_{\max}\right)$,
$\sum_{u\in\mathcal{U}_{K}}\mathbb{P}\left\{B_{u}\right\}=\mathbb{P}\{S\leq K\}=1-\mathbb{P}\{S\geq K+1\}\geq 1-C(z)e^{-(K+1)\ln z}.$ (9)
Eq. 9 establishes that the stationary probability of the truncated state space converges to 1 exponentially fast in $K$. As an illustrative example of a light-traffic two-server system with $\rho_{1}=\rho_{2}=1/2$ and $z=3/2$, we find that $C(z)=4$, and hence $\sum_{u\in\mathcal{U}_{K}}\mathbb{P}\{B_{u}\}\geq 1-4e^{-(K+1)\ln(3/2)}=1-4\left(\tfrac{2}{3}\right)^{K+1}.$ In particular, when taking $K=10$, the stationary probability of the truncated state space $\mathcal{U}_{K}$ exceeds $0.95$. This example demonstrates that the truncated hyperlattice model introduced in Section 3.3.1 provides a valid and accurate approximation in the light-traffic regime.
Join-the-Idle-Queue policy.
Consider a Join-the-Idle-Queue policy in which each arrival, if possible, is routed uniformly at random to one of the idle servers. When all servers are either busy or all idle, the policy reduces to a static random-routing policy $\eta$. The dynamics under this policy are more intricate, as they are state-dependent and must distinguish whether each queue is empty or not. However, this system can be coupled with the one operating under the pure static random-routing policy $\eta$. In particular, let $S^{\ast}(t)$ denote the total number-in-system process (i.e., a birth-death process) under the Join-the-Idle-Queue policy, and let $S(t)$ denote the corresponding coupled process under the pure random-routing policy $\eta$. One can show that $S^{\ast}(t)\leq S(t)$ almost surely for any time $t>0$, since both processes have identical arrival (birth) rates, but under the Join-the-Idle-Queue policy the total service (death) rate is never smaller than under static random routing, as it ensures that at least as many servers remain busy at all times. Therefore, $S^{\ast}$ is stochastically dominated by $S$, implying that the stationary probabilities satisfy $\mathbb{P}\{S^{\ast}<K\}\leq\mathbb{P}\{S<K\}$ (see, e.g., classical results on stochastic comparison of birth-death processes in Van Doorn (2012)). That is, the tail probability bound in Eq. (9) also applies to $S^{\ast}$ under the Join-the-Idle-Queue policy, and the truncated hyperlattice model remains a valid approximation under the light-traffic regime.
3.3.3 Generalization of the hypercube model
The hyperlattice queueing model serves as a natural generalization of Larson’s hypercube model, introduced initially in Larson (1974). Specifically, when the entire region is treated as a single overlapping service region for all servers, i.e., $\mathcal{E}=\{\mathcal{I}\}$, the hyperlattice model reduces to the hypercube model under a given dispatch policy. In the hypercube model, the dispatch policy $\eta$ follows a “fixed preferences” rule, where the dispatcher always selects the highest-ranked available unit based on a predefined preference list, such as the closest-distance dispatching policy. Additionally, the basic hypercube model in Larson (1974) is formulated with each server as an Erlang loss queue with zero capacity. This approach is analogous to our truncated hyperlattice model, except that we extend the loss framework to the entire system rather than individual servers.
Beyond these similarities, our hyperlattice queueing model extends the hypercube model in two fundamental ways. First, it differentiates between overlapping and primary service regions, whereas the hypercube model assumes a single, fully overlapping region. Second, it enables a more flexible, state-dependent dispatch policy that accounts for overlapping services. In the experiments in EC2, we demonstrate that the hyperlattice queueing model provides a more accurate approximation than the hypercube model under complex queueing dynamics with overlapping services.
@@ 3.4 Measures of performance
In this section, we explicate two key performance measures for a service system with overlapping service regions, using the proposed hyperlattice queueing model as a means of analysis.
@@ 3.4.1 Individual workloads
The individual workloads can be further used to calculate various types of system-wide workload imbalance defined in Larson (1974).
@@ 3.4.2 Fraction of dispatches
In addition, we use $\rho_{i,i}$ to represent the fraction of dispatches that send the server $i$ to its primary service region $i$, which is consistently equal to $\lambda_{i}/\lambda$ since this region can exclusively be handled by server $i$. We note that $\rho_{i,e}=0$ if $i\notin e$ and $\sum_{i,e}\rho_{i,e}+\sum_{i}\rho_{i,i}=1$. Additionally, in practice, we consider $u\in\mathcal{U}_{K}$ rather than $\mathcal{U}$ in (11) as we only calculate the steady-state probabilities for the first $U_{K}$ states according to Section 3.3.
3.4.3 Travel time estimation
The proportion of dispatches can be employed to derive expressions for travel time metrics. These metrics can be obtained similarly according to Larson (1974). For any server $i\in\mathcal{I}$ and one of its primary overlapping service regions with server set $e\in\mathcal{E}$, let $t_{i,e}$ denote the mean travel time for sending server $i$ to the service region $\mathcal{O}_{e}$, and let $t_{i,i}$ denote the mean travel time for sending server $i$ to its primary service region. The unconditional mean travel time, denoted as $\bar{\tau}$, can be estimated by
${\bar{\tau}}=\sum_{i\in\mathcal{I}}\left(\sum_{e\in\mathcal{E}}\rho_{i,e}t_{i,e}+\rho_{i,i}t_{i,i}\right).$ (12)
Similarly, the mean travel time for each individual server $i\in\mathcal{I}$, denoted as $\tau_{i}$, can be estimated by
${\tau}_{i}=\frac{\sum_{e\in\mathcal{E}}\rho_{i,e}t_{i,e}+\rho_{i,i}t_{i,i}}{\sum_{e\in\mathcal{E}}\rho_{i,e}+\rho_{i,i}},$ (13)
and the mean travel time for each overlapping service region $e\in\mathcal{E}$, denoted as $\tau_{e}$, is given by
${\tau}_{e}=\frac{\sum_{i\in\mathcal{I}}\rho_{i,e}t_{i,e}}{\sum_{i\in\mathcal{I}}\rho_{i,e}}.$ (14)
However, it is important to note that this approach assumes that queueing delay is the predominant factor influencing the server’s response time for each incident. This assumption has also been discussed in the hypercube queueing model by Larson (1974). Moreover, in this metric, travel times are not explicitly embedded within the model formulation. Instead, following Larson’s hypercube approach, travel times are treated as exogenous factors and assessed through post-hoc analyses.